Grade 12th passMagical Mathematics[Interesting Approach]

Evaluateintegrate (1/a+bcosx).dx limit 0 to p

Miss Pooja
12 Years agoGrade 12th pass

1 Answer

Arun Kumar
12 Years ago
Hello Student,
integrate dx/(a+bcosx)
put cosx=(1-tan^2(x/2))/(1+tan^(x/2))
integrate sec^2(x/2) dx/a(1+tan^2(x/2)+b(1-tan^2(x/2))
=integrate sec^2(x/2)dx/(a+b)+(a-b)tan^2(x/2)
put tan(x/2)=t
diff w. r. to x we get
sec^2(x/2).1/2 dx=dt
=integrate 2dt/(a+b)+(a-b)t^2
=2/sqrt(a+b) arctan( sqrt(a-b)t/sqrt(a+b))
=2/sqrt(a+b) arctan(sqrt(a-b)tan(x/2)/sqrt(a+b))
Thanks & Regards
Arun Kumar
Btech, IIT Delhi
Askiitians Faculty